OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -1 and a y-intercept of 2? y = 2x - 2 y = - 2x + 2 OR y = 2x + 2

OpenStudy (anonymous):

 look if the line is ascendent, x may be positive. When x=0, y=2 and when y=0, x=-1 y= 2x +2

OpenStudy (anonymous):

ah! you can use too: y-y0 = m(x-x0) at (0, 2) y-2 = m (x-0) y=mx +2 [1] m is the slope: so taking two points (0 ,2) and ( -1 , 0) m = (2 - 0)/( 0 -(-1)) = 2 replacing in [1] y= 2x +2